a. r =

=

=

=

=

=

r =


b. There is a _____________________between the self-concept skill and leadership skill of the students.


x y xy x2 y2

9.5 9.2

9.2 8.8

6.3 7.3

4.1 3.4

5.4 6.0

8.3 7.8

7.8 8.8

6.8 7.0

5.6 6.5

7.1 8.3

"\sum"= "\sum"= "\sum"= "\sum"= "\sum"=


c)


= a + bx a= ( "\sum" y)("\sum" x2) - ("\sum" x)("\sum" xy) / n("\sum" x2) -"\sum" x2

= = = a=

b= ("\sum" xy) - ("\sum" x) ("\sum" y) / n("\sum"x2 ) -("\sum" x)2


d)= _____________

= ____________(1.5)_

= ____________



1
Expert's answer
2022-06-13T10:47:06-0400

In order to compute the regression coefficients, the following table needs to be used:




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & X & Y & XY & X^2 & Y^2 \\ \hline
 & 9.5 & 9.2 & 87.4 & 90.25 & 84.64 \\
 \hdashline
 & 9.2 & 8.8 & 80.96 & 84.64 & 77.44 \\
 \hdashline
 & 6.3 & 7.3 & 45.99 & 39.69 & 53.29 \\
 \hdashline
 & 4.1 & 3.4 & 13.94 & 16.81 & 11.56 \\
 \hdashline
 & 5.4 & 6.0 & 32.4 & 29.16 & 36 \\
 \hdashline
 & 8.3 & 7.8 & 64.74 & 68.89 & 60.84 \\
 \hdashline
 & 7.8 & 8.8 & 68.64 & 60.84 & 77.44 \\
 \hdashline
 & 6.8 & 7.0 & 47.6 & 46.24 & 49 \\
 \hdashline
 & 5.6 & 6.5 & 36.4 & 31.36 & 42.25 \\
 \hdashline
 & 7.1 & 8.3 & 58.93 & 50.41 & 68.89 \\
 \hdashline
Sum= & 70.1 & 73.1 & 537 & 518.29 & 561.35 \\
 \hdashline
\end{array}"





"\bar{X}=\dfrac{1}{n}\sum _{i}X_i=\dfrac{70.1}{10}=7.01"





"\bar{Y}=\dfrac{1}{n}\sum _{i}Y_i=\dfrac{73.1}{10}=7.31"





"SS_{XX}=\sum_iX_i^2-\dfrac{1}{n}(\sum _{i}X_i)^2""=518.29-\dfrac{70.1^2}{10}=26.889"





"SS_{YY}=\sum_iY_i^2-\dfrac{1}{n}(\sum _{i}Y_i)^2""=561.35-\dfrac{(73.1)^2}{10}=26.989"






"SS_{XY}=\sum_iX_iY_i-\dfrac{1}{n}(\sum _{i}X_i)(\sum _{i}Y_i)""=537-\dfrac{70.1(73.1)}{10}=24.569"


a.




"r=\dfrac{SS_{XY}}{\sqrt{SS_{XX}SS_{YY}}}=\dfrac{24.569}{\sqrt{26.889(26.989)}}"





"=0.912025"


b. There is a strong positive correlation between the self-concept skill and leadership skill of the students.


c.




"m=slope=\dfrac{SS_{XY}}{SS_{XX}}=\dfrac{24.569}{26.889}=0.9137""n=\bar{Y}-m\bar{X}=7.31-\dfrac{24.569}{26.889}(7.01)=0.9048"



The regression equation is:



"y=0.9048+0.9137x"

d.

"x=1.5"




"y=0.9048+0.9137(1.5)=2.275"